OpenStudy (anonymous):

If the lower base of an isosceles trapezoid is three times as longs as the upper base and one of its legs measures 10cm, what is the lenght of the lower base if the perimeter of the trapezoid is 36cm?

OpenStudy (anonymous):

hey....this is a very easy question cos perimeter= x+3x+(2*10) = 36 so. 4x=16 and x=4cm so lower base = 3x = 3*4 = 12 so length of lower base is 12 cm

OpenStudy (anonymous):

sorry im not good at math. Thanks for the help

OpenStudy (anonymous):

@ankitisphysicist where did you get 2?

OpenStudy (anonymous):

it is in question that trapezoid is isosceles that means both its legs are equal ..so i have taken on leg to be y therefore sum of two legs will be 2y or you can say 2*10

OpenStudy (anonymous):

@ankitisphysicist thanks again !

OpenStudy (anonymous):

most welcome